# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

### v2

@@ Results
Results from an explicit simulation of this simplified model are shown in Fig. 3 (dots). A mean-field analytic approximation can be found by fixing the source unit in the center of a domain, assigning a diffusivity of $D=2D_{n}$ to the remaining units and solving for the spatial concentration field $h(r,t)$, which defines the mean concentration per network unit located at distance $r$ from the origin. The average of this field within the narrow contact zone is defined as $h_{c}(t)$. The time evolution of the concentration fields can be expressed as follows (details in SI Appendix):
$\frac{dh(r,t)}{dt}$ $=D\nabla^{2}h(r,t)-k_{d}h(r,t),\;\;\text{for}\ b<r<R$ (2a)
$\frac{dh_{c}(t)}{dt}$ $=k_{u}\left[h_{0}\ell_{0}-h_{c}(t)\right]-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (2b)
@@
where $k_{u}\left[h_{0}\ell_{0}-h_{c}(t)\right]$ represents the injection of new material into the system via fusion with the source, $I=-4\pi b^{2}\rho D\left.\frac{dh(r,t)}{dr}\right|_{b}$ is the current of material leaving the contact zone, $\rho=(N_{0}-\langle n\rangle)/V$ is the density of network units, $V=4/3\pi(R^{3}-a^{3})$ is the domain volume, and $v_{c}=4/3\pi(b^{3}-a^{3})$ is the contact zone volume. Interactions between non-source units do not alter the mean-field concentrations, and the overall material in the system can only increase upon encounters with the source.
@@
Figure 3: Material delivered from a source unit in a system of fragmented clusters. Solid lines show mean-field solutions (see Eq. 3) with unit cluster size $\langle n\rangle=1$, for different fusion rates $k_{u}$. Colored dots show explicit simulation results for a simplified system of interacting spheres with uniform size (inset). Dashed black lines show the limiting behavior for fusion-limited and encounter-limited regimes.
@@
We compute the steady-state solution of Eq. 2 and integrate $h(r)$ over space to find the total amount of material ($S$) in the system, excluding the source unit (details in SI Appendix):
$S=$ $(\langle n\rangle-1)+\rho h_{c}(v_{c}+z),$ (3a)
$h_{c}=$ $\frac{1}{1+k_{d}/k_{u}+k_{d}z/(k_{u}v_{c})},$ (3b)
$\frac{z}{4\pi b\lambda}=$ $\frac{(\lambda-b)(R+\lambda)+(\lambda+b)(R-\lambda)e^{2(R-b)/\lambda}}{(R+\lambda)+(R-\lambda)e^{2(R-b)/\lambda}},$ (3c)
@@
where $\lambda=\sqrt{D/k_{d}}$ is the diffusive lengthscale for spreading through the population of clusters. Here, the first term in $S$ represents material in the cluster containing the source unit, the term $\rho h_{c}v_{c}$ is the total material within the contact layer (defined by a balance of fusion, decay, and diffusive escape), and $z$ describes the additional volume over which the material has spread beyond the contact zone.
@@
The total material in the system exhibits distinct scaling behaviors (Fig. 3, dashed) depending on the relative timescales for decay, encounter ($\tau_{\text{enc}}=(4\pi Db\rho)^{-1}$), and fusion ($\tau_{u}=(k_{u}\rho v_{c})^{-1}$). In keeping with standard results for diffusion to a partially reactive target [73], the overall interaction time is given by the sum of waiting times for the two-step process: $\tau_{\text{int}}=\tau_{\text{enc}}+\tau_{u}$. In the limit of fast decay ($\tau_{d}\ll\tau_{\text{int}}$), the steady-state material in a social network of unit-size clusters approaches $S\rightarrow\rho v_{c}h_{c}=\rho v_{c}k_{u}/k_{d}$. In this limit, the total amount delivered is set by the number of units in the contact zone ($\rho v_{c}$) and how often they fuse with the source during the decay time ($k_{u}/k_{d}$), with no spatial spread.
@@
Another limit arises when diffusion of clusters is fast relative to both decay and fusion ($\tau_{\text{enc}}\ll\tau_{d},\tau_{u}$). This yields the same scaling for the total network content $S\rightarrow\rho v_{c}k_{u}/k_{d}$, with $h_{c}\rightarrow k_{u}v_{c}/4\pi Db$. The amount of material in the contact layer ($h_{c}$) is set by a balance between injection through fusion (at rate $k_{u}\rho v_{c}$) and escape through the diffusive arrival of fresh units that dilute the local concentration ($4\pi Db\rho$). The additional volume $z\rightarrow 4\pi Db/k_{d}$ over which the material spreads depends on the balance between diffusive encounters and decay. Overall, the total amount of material in the network is fusion-limited as the rapid encounters quickly homogenize the individual units within the diffusive range of the source.
@@
Fig. 3 shows the full solution of the mean-field model (Eq. 3) for spreading on a fragmented network, with variable fusion rate $k_{u}$ and a unit cluster size $\left<n\right>=1$. The analytic calculations accurately reproduce simulations with diffusive spheres that exchange material via transient fusion (dots in Fig. 3, details in SI Appendix).
Figure 4: Spreading on simulated networks exhibits a transition between the static network and social network regime. The total material is plotted as a function of the decay time for different fusion rate constants $k_{u1}$ (solid lines). The analytic approximation (Eq. 4, dashed curves) encompases both the fractal continuum limit (for large clusters or fast decay) and the social network limit (for small clusters and slow decay). Plateaus at intermediate decay times correspond to filling of an individual cluster before interactions can occur. Insets show example simulation snapshots for the parameters indicated by the gray arrows. Parameters $N_{0}=250,\ell_{0}=0.5,R=5,k_{u2}=\frac{1}{3}k_{u1},k_{f}=1$ are used in both (a) and (b), with the particle diffusivity reduced from $D_{p}=4800$ in (a) to $D_{p}=48$ in (b).
@@
The model for spreading through a social network of interacting fragments can be expanded to approximate a regime with larger clusters. We use simulations of dynamic networks with different structures [57] to extract the effective parameters (mean cluster size $\left<n\right$>, effective unit diffusivity $D_{n}$, steric radius $a_{n}$, contact volume $v_{c}$, and effective local fusion rate $k_{u}$) for the simplified social network model (details in SI Appendix).
@@
When individual clusters are homogeneously filled upon each interaction ($\tau_{c}<\left\{\tau_{d},\tau_{\text{int}},\tau_{f}\right\}$), the solution in Eq. 3 can be used directly with the appropriate values of the parameters $a,v_{c},k_{u}$. However, for slowly diffusing particles or very large clusters ($\tau_{c}>\tau_{d}$), the source cluster may be only partly full of material. The average concentration in the source cluster, $h_{0}^{(d)}$, is then determined by the static network solution (Eq. 1) within the cluster: $\ell_{0}h_{0}^{(d)}=(S_{d}(\sqrt{D_{p}/k_{d}})+1)/\langle n\rangle$. When another cluster fuses with the source cluster, fission may terminate the encounter before filling is complete ($\tau_{f}<\tau_{c}$). The amount transferred should then be scaled by the fraction of the new cluster that is filled prior to fission: $f=(S_{d}(\sqrt{D_{p}/k_{f}})+1)/\langle n\rangle$. These two corrections modify the dynamic equation for the concentration in the contact zone (Eq. 2b) and its solution (Eq. 3a,b) as follows:
$\frac{dh_{c}(t)}{dt}$ $=k_{u}\left[\ell_{0}h_{0}^{(d)}-h_{c}(t)\right]f-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (4a)
$S$ $=S_{d}\left(\sqrt{D_{p}/k_{d}}\right)+\rho h_{c}(v_{c}+z),$ (4b)
$h_{c}$ $=\frac{\ell_{0}h_{0}^{(d)}}{1+k_{d}/(fk_{u})+k_{d}z/(fk_{u}v_{c})},$ (4c)
@@
where the first term in the updated $S$ accounts for material in the source cluster and the second term represents material in the rest of the network. When clusters are large and interactions are infrequent, the above solution reduces to the static network limit of Eq. 1. When the clusters are small, the source cluster is fully filled ($h_{0}^{(d)}\rightarrow h_{0}$) and so is each cluster that interacts with it ($f\rightarrow 1$). The general solution then approaches the fragmented social network limit of Eq. 3.
@@
In Fig. 4, we show the spreading of material through simulated networks with increasing connectivity, spanning across the different regimes. In the limit of rapid decay (small $\tau_{\textrm{d}}$), material spreads within a single cluster of dimension $d$, with the amount delivered scaling as $S\sim\tau_{d}^{d/2}$ When particle diffusivity is fast and clusters are small, the source cluster fills faster than interactions can occur and the amount of material plateaus at the size of a single cluster, in the regime of $\tau_{c}<\tau_{d}<\tau_{\text{int}}$. For longer decay times, interaction between distinct clusters dominates the spread, and the total material scales linearly as $S\sim\tau_{d}$. When particle diffusion is slow (Fig. 4b), the cluster containing the source unit is able to offload material through interactions even before it is fully filled ($\tau_{\text{int}}<\tau_{c}$), and the plateau region narrows or disappears.
@@
The simplified analytic model (Eq. 4) approximately matches simulations over a broad range of network connectivities and timescales (Fig. 4). However, some discrepancy occurs for networks that are close to the percolation transition (light green curve), particularly when the particle diffusivity is low. In this case, the approximation systematically overestimates material spreading through the network. This may be the result of transient fissions within a single cluster (which could temporarily hinder material spreading [21]), or steric inaccessibility of nodes buried within a cluster, neither of which are accounted for in our analytic models. Additionally, the percolation transition corresponds to a broad variability in cluster sizes [13], resulting in the source unit occasionally being trapped in very small clusters that limit material delivery.
@@
The quantitative description of steady-state network filling (Eq. 4) can be generalized to other measures of material spread on a dynamic network. In particular, the decay time $\tau_{d}$ sets the timescale over which the spreading is assessed. The total amount of material delivered into the network over time is also well-approximated by the physical and social network models presented here (see Supplemental Fig. S1 in SI Appendix).
@@
To illustrate the utility of our model in providing directly testable predictions for mitochondrial mixing, we analyze time-resolved images of fluorescently tagged mitochondria in three human cell lines: SH-SY5Y, IMR90, and U2OS. Specifically, we segment, skeletonize, and track the mitochondria of each cell in our data set (Figure 5a and Supplemental Video 2), extracting structural metrics (degree distribution, cluster size, fractal dimension) and dynamic parameters (fission rate and mitochondrial diffusivity). These values (provided in Supplemental Material) are used to parameterize our analytic model, which provides a quantitative estimate of mitochondrial material spreading. We estimate the time required for material to fill half the network ($t_{1/2}$) as the value of $\tau_{d}$ corresponding to $S$ reaching half its maximal value. This metric provides a prediction for how quickly particles of different types should spread through most of the mitochondrial network.
@@
Figure 5b shows our predictions of spreading time on prototypical ‘average cells’, with individual parameters averaged across all the cells of that type. The spreading time is plotted as a function of the material diffusivity, $D_{p}$, in real units. For typical mitochondrial matrix proteins with $D_{p}=20\mu\text{m}^{2}\text{/s}$ [66], we expect a spreading time of $\sim 4$ min for SH-SY5Y, and $\sim 0.5$ hr for IMR90 and U2OS cells. Notably, the different cell types exhibit distinct behaviors. For IMR90 and U2OS cells, the spreading time asymptotes for particle diffusivities above roughly $1\mu\text{m}^{2}/\text{s}$. By contrast, in the typical SH-SY5Y cell, spreading times continue to decrease with higher material diffusivity. These results arise primarily from differences in the network architecture: SH-SY5Y mitochondria tend to be hyperfused, while IMR90 and U2OS exhibit more disconnected networks, so that dispersion becomes limited by interactions between distinct mitochondrial clusters rather than diffusion within connected mitochondria. Overall, our results indicate that cultured human cell lines can vary substantially in their mitochondrial architecture and mixing dynamics.
@@
It should be noted, however, that there is extensive variation in the mitochondrial structure and associated network parameters between individual cells of the same type (dashed lines in Fig. 5b). To further explore the dependence of spreading rates on network connectivity, we reclassified all cells in our data set as ‘disconnected’ or ‘hyperfused’ based on the mean cluster size as a fraction of the total mitochondrial population. Figure 5c shows the distribution of cells by mean cluster size fraction and the classification threshold obtained via a k-means clustering algorthim. All cells with mean cluster size fraction below the threshold were classified as disconnected, with the remainder considered hyperfused. Figure 5d demonstrates the difference in spreading behavior for these two cell classes, with hyperfused cells exhibiting strong dependence on material diffusivity across the entire relevant range. Spreading on disconnected networks plateaus for particles with diffusivities above $D_{p}\approx 1\mu\text{m}^{2}\text{/s}$. We note that for slowly diffusing material, disconnected networks can actually increase the spreading rate. This is due to the mobility of mitochondrial clusters, which allows each unit to interact with more neighbors than they could contact in the static low-dimensional hyperfused networks ($1\leq d\lesssim 2$).
@@
The mitochondrial networks for different cell types exhibit multiple morphological differences. In particular, the total network size in U2OS cells is substantially larger than in the other two cell types (Fig. 5a; see Supplemental Material). However, because the mean cluster size fraction is similar in both IMR90 and U2OS cells, the predicted filling half-time is similar as well. For systems that fall into the disconnected social network regime, it is the relative size of the individual mitochondrial clusters compared to the full network that determines filling times.
@@
The distinction in predicted spreading rates between different cell types can be attributed to their different propensities for mitochondrial connectivity. Many (though not all) SH-SY5Y cells exhibit hyperfused networks, while IMR90 and U2OS cells fall primarily in the disconnected category. SH-SY5Y cells thus serve as examples of the stationary physical network regime, where spreading is limited by material transport along a fractal-dimensional continuum. The other two cell types have mitochondria that behave akin to a social network, with interactions between clusters limiting the spread of rapidly diffusing particles. Notably, for both IMR90 and U2OS mitochondrial networks, the spreading at high $D_{p}$ is limited by mitochondrial mobility rather than the timescale of local fusion or fission (see Supplemental Appendix). We note that the balance of fusion to fission rates, which sets network connectivity, plays an important role in dictating material spread. However, our analysis indicates that the interaction rate is limited by mitochondria moving through space to encounter new fusion partners rather than by the waiting time for fusion between two nearby mitochondria ($\tau_{\text{enc}}\gg\tau_{u}$). Thus, the model predicts that proportional increases in fusion and fission rates should have little effect on mitochondrial mixing.

### v3

@@ Results
Results from an explicit simulation of this simplified model are shown in Fig. 3. A mean-field analytic approximation can be found by fixing the source unit in the center of a domain, assigning a relative diffusivity of $D=2D_{n}$ to the remaining units and solving for the spatial field $h(r,t)$, which defines the material per network unit located at distance $r$ from the origin. The average of this field within the narrow contact zone is defined as $h_{c}(t)$. The time evolution of the concentration fields can be expressed as follows (details in SI Appendix): {subequations} {align} ∂h(r,t)∂t = D∇^2 h(r,t)-k_dh(r,t),    for b¡r¡R
@@
dhc(t)dt = k_u[h_0ℓ_0-h_c(t)] - Iρvc -k_dh_c(t), where $k_{u}\left[h_{0}\ell_{0}-h_{c}(t)\right]$ represents the injection of new material into the system via fusion with the source, $I=-4\pi b^{2}\rho D\left.\frac{\partial h(r,t)}{\partial r}\right|_{b}$ is the current of material leaving the contact zone, $\rho=(N_{0}-\langle n\rangle)/V$ is the density of network units, $V=\frac{4}{3}\pi(R^{3}-a^{3})$ is the domain volume, and $v_{c}=\frac{4}{3}\pi(b^{3}-a^{3})$ is the contact zone volume. Interactions between non-source units do not alter the mean-field concentrations, and the overall material in the system can only increase upon encounters with the source.
@@
Figure 3: Steady-state material content supplied by a source unit in a system of fragmented clusters. Solid lines show mean-field solutions (Eq. 3) with unit cluster size $\langle n\rangle=1$, for different fusion rates $k_{u}$. Colored dots show explicit simulation results for a simplified system of interacting spheres with uniform size (inset). Dashed black lines show the limiting behavior for fusion-limited and encounter-limited regimes.
@@
We compute the steady-state solution of Eq. Fragmented network regime: spreading through transient interactions and integrate $h(r)$ over space to find the total amount of material ($S$) in the system, excluding the source unit (details in SI Appendix): {subequations} {align} S = h_0ℓ_0(⟨n ⟩- 1) + h_c ρ(v_c+z),
h_c = h001+kd/ku+kdz/(kuvc),
@@
z4πbλ = (λ-b)(R+λ)+(λ+b)(R-λ)e2(R-b)/λ(R+λ) + (R-λ)e2(R-b)/λ, where $\lambda=\sqrt{D/k_{d}}$ is the diffusive lengthscale for spreading through the population of clusters. Here, the first term in $S$ represents material in the cluster containing the source unit, the term $h_{c}\rho v_{c}$ is the total material within the contact layer (defined by a balance of fusion, decay, and diffusive escape), and $z$ describes the additional volume over which the material has spread beyond the contact zone.
@@
The total material in the system exhibits distinct scaling behaviors (Fig. 3) depending on the relative timescales for decay, encounter ($\tau_{\text{enc}}=(4\pi Db\rho)^{-1}$), and fusion ($\tau_{u}=(k_{u}\rho v_{c})^{-1}$). In keeping with standard results for diffusion to a partially reactive target [99], the overall interaction time is given by the sum of waiting times for the two-step process: $\tau_{\text{int}}=\tau_{\text{enc}}+\tau_{u}$. In the limit of fast decay ($\tau_{d}\ll\tau_{\text{int}}$), the steady-state material in a social network of unit-size clusters approaches $S\rightarrow\rho v_{c}h_{c}=\rho v_{c}k_{u}/k_{d}$. In this limit, the steady-state content is set by the number of units in the contact zone ($\rho v_{c}$) and how often they fuse with the source during the decay time ($k_{u}/k_{d}$), with no spatial spread.
@@
Another limit arises when diffusion of clusters is fast relative to both decay and fusion ($\tau_{\text{enc}}\ll\tau_{d},\tau_{u}$). This yields the same scaling for the steady-state content $S\rightarrow\rho v_{c}k_{u}/k_{d}$, with $h_{c}\rightarrow k_{u}v_{c}/(4\pi Db)$. The amount of material in the contact layer ($h_{c}$) is set by a balance between injection through fusion (at rate $k_{u}\rho v_{c}$) and escape through the diffusive arrival of fresh units that dilute the local concentration ($4\pi Db\rho$). The additional volume $z\rightarrow 4\pi Db/k_{d}$ over which the material spreads depends on the balance between diffusive encounters and decay. Overall, the total amount of material in the network is fusion-limited as the rapid encounters quickly homogenize the individual units within the diffusive range of the source. We note that this limit is analogous to the homogeneous-mixing limit described for epidemic spread among rapidly diffusing mobile agents [4].
@@
Fig. 3 shows the full solution of the mean-field model (Eq. 3) for spreading on a fragmented network, with variable fusion rate $k_{u}$ and a unit cluster size $\left<n\right>=1$. The analytic calculations accurately reproduce simulations with diffusive spheres that exchange material via transient fusion (details in SI Appendix).
For simplicity, our model assumes effectively diffusive motion of mitochondrial units. However, individual mitochondria in both plant and animal cells can also engage in directed motor-driven runs [35, 59, 100]. Such motion increases the effective diffusivity of the mitochondria over timescales encompassing multiple runs, and concomitantly increases the encounter rate between particles [4, 101]. The effect of processive runs on network filling in the fragmented regime is shown in Supplemental Fig. S1 in the SI Appendix. As reported previously, long runs allow for more rapid mixing of network contents [59], with diminishing effect once the run-length becomes comparable to the domain size [101, 4].
Figure 4: Spreading on simulated networks exhibits a transition between the static network and social network regime. The steady-state material content is plotted as a function of the decay time for different values of the fusion rate constant $k_{u1}$ (solid lines), which tunes between fragmented networks (blue), networks at the percolation transition (light green) and highly fused networks (yellow). The analytic approximation (Eq. Dynamic networks with large interacting clusters, dashed curves) encompases both the fractal continuum limit (for large clusters or fast decay) and the social network limit (for small clusters and slow decay). Plateaus at intermediate decay times correspond to filling of an individual cluster before interactions can occur. Insets show example simulation snapshots for the parameters indicated by the gray arrows. Parameters $N_{0}=250,\ell_{0}=0.5,R=5,k_{u2}=\frac{1}{3}k_{u1},k_{f}=1$ are used in both (a) and (b), with the particle diffusivity set to $D_{p}=4800$ in (a) and $D_{p}=48$ in (b).
@@
The model for spreading through a social network of interacting fragments can be expanded to approximate a regime with larger clusters. We use simulations of dynamic networks with different structures [67] to extract the effective parameters (mean cluster size $\left<n\right$>, effective unit diffusivity $D_{n}$, steric radius $a$, contact volume $v_{c}$, and effective local fusion rate $k_{u}$) for the simplified social network model (details in SI Appendix). The encounter time between clusters decreases with cluster size but remains well-approximated by the expression for diffusion-limited reactions of spherical particles, $\tau_{\text{enc}}=1/(4\pi Db\rho)$ [102] (Supplemental Fig. S2 in SI Appendix).
@@
When individual clusters are homogeneously filled upon each interaction ($\tau_{c}<\left\{\tau_{d},\tau_{\text{int}},\tau_{f}\right\}$), the solution in Eq. 3 can be used directly with the appropriate values of the parameters $a,v_{c},k_{u}$. However, for slowly diffusing particles or very large clusters ($\tau_{c}>\tau_{d}$), the source cluster may be only partly full of material. The average concentration in the source cluster, $h_{0}^{(d)}$, is then determined by the static network solution (Eq. 1) within the cluster: $\ell_{0}h_{0}^{(d)}=(S_{d}(\sqrt{D_{p}/k_{d}})+h_{0}\ell_{0})/\langle n\rangle$. When another cluster fuses with the source cluster, fission may terminate the encounter before filling is complete ($\tau_{f}<\tau_{c}$). The amount transferred should then be scaled by the fraction of the new cluster that is filled prior to fission: $f=(S_{d}(\sqrt{D_{p}/k_{f}})+h_{0}\ell_{0})/\langle n\rangle$. These two corrections modify the dynamic equation for the material in the contact zone (Eq. Fragmented network regime: spreading through transient interactionsb) and its solution (Eq. 3a,b) as follows: {subequations} {align} dhc(t)dt = k_u [ℓ_0 h_0^(d) -h_c(t) ]f -Iρvc-k_dh_c(t),
S = S_d(D_p/k_d) + ρh_c(v_c+z),
@@
h_c = ℓ0h0(d)1+kd/(f ku)+kdz/(fkuvc), where the first term in the updated $S$ accounts for material in the source cluster and the second term represents material in the rest of the network. When clusters are large and interactions are infrequent, the above solution reduces to the static network limit of Eq. 1. When the clusters are small, the source cluster is fully filled ($h_{0}^{(d)}\rightarrow h_{0}$) and so is each cluster that interacts with it ($f\rightarrow 1$). The general solution then approaches the fragmented social network limit of Eq. 3.
@@
In Fig. 4, we show the spreading of material through simulated networks with increasing connectivity, spanning across the different regimes. In the limit of rapid decay ($\tau_{\textrm{d}}<\tau_{\textrm{c}}$), material spreads within a single cluster of dimension $d$, with the steady-state content scaling as $S\sim\tau_{d}^{d/2}$. When particle diffusivity is fast and clusters are small, the source cluster fills faster than interactions can occur and the amount of material plateaus at the size of a single cluster, in the regime of $\tau_{c}<\tau_{d}<\tau_{\text{int}}$. This plateau implies that network filling is stalled over a certain timescale, consistent with past results on dynamic lattice networks below the percolation transition, when edge fluctuations are slow [22, 24]. The plateau is expected to disappear when edges turn on and off rapidly, restoring network connectivity over longer timescales. In our system, this corresponds to longer decay times, for which interactions between distinct clusters dominate the spread. In this long-time regime, spreading through the population is three-dimensional and the total network content scales linearly as $S\sim\tau_{d}$.
When particle diffusion is slow (Fig. 4b), the cluster containing the source unit is able to offload material through interactions even before it is fully filled, and the plateau region narrows. If the rate of novel encounters is faster than the intra-cluster spreading timescale ($\tau_{\text{int}}<\tau_{c}$), the plateau disappears, as observed in lattice networks with rapid bond fluctuations [22, 24].
@@
The simplified analytic model (Eq. Dynamic networks with large interacting clusters) approximately matches simulations over a broad range of network connectivities and timescales (Fig. 4). However, some discrepancy occurs for networks that are close to the percolation transition, particularly when the particle diffusivity is low. In this case, the approximation systematically overestimates material spreading through the network. This may be the result of transient fissions within a single cluster (which could temporarily hinder material spreading [24]), or steric inaccessibility of nodes buried within a cluster, neither of which are accounted for in our analytic models. Additionally, the percolation transition corresponds to a broad variability in cluster sizes [15], resulting in the source unit occasionally being trapped in very small clusters that limit material delivery.
We note that in the limit of arbitrarily slow particle diffusivity or fragmented networks, mitochodrial content would remain limited to units which have come in direct contact with the source unit over a timescale of $\tau_{d}$. Steady-state filling is then determined by the cumulative degree of the temporal network describing all novel fusions made by a mitochondrial unit. This degree is plotted as a function of time in Supplemental Fig. S2. Notably, the novel contact degree is highest for moderate cluster sizes, which enable rapid rearrangement of the network structure, as highlighted in prior work [67]. The ability of fragmented social networks of mobile mitochondria to efficiently share slowly-moving genomic elements has also been pointed out in the context of plant mitochondrial dynamics [59, 49].
@@
The quantitative description of steady-state network filling (Eq. Dynamic networks with large interacting clusters) can be generalized to other measures of material spread on a dynamic network. In particular, the decay time $\tau_{d}$ sets the timescale over which the spreading is assessed. The total amount of material delivered into the network over time is also well-approximated by the physical and social network models presented here (see Supplemental Fig. S3 in SI Appendix).
@@
To illustrate the utility of our model in providing directly testable predictions for mitochondrial mixing, we analyze time-resolved images of fluorescently tagged mitochondria in three human cell lines: SH-SY5Y, IMR90, and U2OS. Specifically, we segment, skeletonize, and track the mitochondria of each cell in our data set (Figure 5a and Supplemental Video 2), extracting structural metrics (degree distribution, cluster size, graph dimension) and dynamic parameters (fission rate and mitochondrial diffusivity). These values (provided in the SI Appendix) are used to parameterize our analytic model, which provides a quantitative estimate of mitochondrial material spreading. We estimate the time required for material to fill half the network ($t_{1/2}$) as the value of $\tau_{d}$ corresponding to $S$ reaching half its maximal value. This metric provides a prediction for how quickly particles of different types should spread through most of the mitochondrial network.
@@
Figure 5b shows our predictions of spreading time on prototypical ‘average cells’, with individual parameters averaged across all the cells of that type. The spreading time is plotted as a function of the material diffusivity, $D_{p}$, in real units. For typical mitochondrial matrix proteins with $D_{p}=20\mu\text{m}^{2}\text{/s}$ [79], we expect a spreading time of $\sim 6$ min for SH-SY5Y, and $\sim 45$ min for IMR90 and U2OS cells. Notably, the different cell types exhibit distinct behaviors. For IMR90 and U2OS cells, the spreading time asymptotes for particle diffusivities above roughly $1\mu\text{m}^{2}/\text{s}$. By contrast, in the typical SH-SY5Y cell, spreading times continue to decrease with higher material diffusivity. These results arise primarily from differences in the network architecture: SH-SY5Y mitochondria tend to be hyperfused, while IMR90 and U2OS exhibit more disconnected networks, so that dispersion becomes limited by interactions between distinct mitochondrial clusters rather than diffusion within connected mitochondria. Overall, our results indicate that cultured human cell lines can vary substantially in their mitochondrial architecture and rate of material transport.
@@
It should be noted, however, that there is extensive variation in the mitochondrial structure and associated network parameters between individual cells of the same type (Fig. 5b). To further explore the dependence of spreading rates on network connectivity, we reclassified all cells in our data set as ‘disconnected’ or ‘hyperfused’ based on the mean cluster size as a fraction of the total mitochondrial population. Figure 5c shows the distribution of cells by mean cluster size fraction and the classification threshold obtained via a k-means clustering algorthim. All cells with mean cluster size fraction below the threshold were classified as disconnected, with the remainder considered hyperfused. Figure 5d demonstrates the difference in spreading behavior for these two cell classes, with hyperfused cells exhibiting strong dependence on material diffusivity across the entire relevant range. Spreading on disconnected networks plateaus for particles with diffusivities above $D_{p}\approx 1\mu\text{m}^{2}\text{/s}$. We note that for slowly diffusing material, disconnected networks can actually increase the spreading rate relative to hyperfused networks. This is due to the mobility of mitochondrial clusters, which allows each unit to interact with more neighbors than they could contact in the static low-dimensional hyperfused networks ($1\leq d\lesssim 2$). The enhanced rate of novel contacts between mitochondrial units in partially disconnected networks is demonstrated in the SI Appendix (Fig. S2d).
@@
The mitochondrial networks for different cell types exhibit multiple morphological differences. In particular, the total network size in U2OS cells is substantially larger than in the other two cell types (Fig. 5a; see SI Appendix). However, because the mean cluster size fraction is similar in both IMR90 and U2OS cells, the predicted filling half-time is similar as well. For systems that fall into the disconnected social network regime, it is the relative size of the individual mitochondrial clusters compared to the full network that determines filling times.
@@
The distinction in predicted spreading rates between different cell types can be attributed to their different propensities for mitochondrial connectivity. Many (though not all) SH-SY5Y cells exhibit hyperfused networks, while IMR90 and U2OS cells fall primarily in the disconnected category. We note that these features may be dependent on the specific growth conditions (described in SI appendix) as well as the cell type. SH-SY5Y cells thus serve as examples of the physical network regime, where spreading is limited by material transport along a fractal-dimensional continuum. The other two cell types have mitochondria that behave akin to a social network, with interactions between clusters limiting the spread of rapidly diffusing particles. Notably, for both IMR90 and U2OS mitochondrial networks, the spreading at high $D_{p}$ is limited by mitochondrial mobility rather than the timescale of local fusion or fission (see SI Appendix for extracted parameters). The balance of fusion to fission rates, which sets network connectivity, plays an important role in dictating material spread. However, our analysis indicates that the interaction rate is limited by mitochondria moving through space to encounter new fusion partners rather than by the waiting time for fusion between two nearby mitochondria ($\tau_{\text{enc}}\gg\tau_{u}$). Thus, the model predicts that proportional increases in fusion and fission rates should have little effect on mitochondrial mixing.
In other organisms and cell types, such as plant cells, mitochondria appear to be significantly more mobile, traversing tens of microns over a few-minute timescale [35]. Their active motion, which relies on processive motor-driven runs, results in rapid exploration throughout the cell, with encounter times apparently much smaller than the fusion timescale $\tau_{\text{enc}}\ll\tau_{u}$ (see SI Appendix). In such a regime, mitochondrial material can be broadly spread out in space while individual units remain heterogeneous, as was observed for the mixing of photoconverted mitochondria in onion bulb epidermal cells [103]. The biological variety of mitochondrial systems thus necessitates consideration of all the distinct timescales discussed in our general model for spreading through temporal networks.